OpenStudy (anonymous):

Evaluate and simplify tan 15º.

OpenStudy (anonymous):

Do not duplicate the question..

OpenStudy (anonymous):

15 = 45-30 so, \[\tan(15^\circ)=\tan(45^\circ-30^\circ)\] you can you the tan (A-B) expansion.

OpenStudy (anonymous):

You are just writing the answer part.. Why you are not writing the solution and calculation part so that we can check where are you doing wrong..??

OpenStudy (anonymous):

2+qrt3 is that correct

OpenStudy (anonymous):

I am again saying you that write the proper solution..

OpenStudy (anonymous):

Then you are guessing the answers??

OpenStudy (anonymous):

im using the answers from the multiple choice

OpenStudy (anonymous):

Let me show you properly: \[\tan(45 - 30) = \frac{\tan(45) - \tan(30)}{1 + \tan(45)\tan(30)}\] \[\implies \frac{1 - \frac{1}{\sqrt{3}}}{1 + 1 \cdot \frac{1}{\sqrt{3}}}\] Can you solve this?

OpenStudy (anonymous):

Just guessing the answer will not help you buddy..

OpenStudy (anonymous):

my answer still isnt coming out right

OpenStudy (anonymous):

I DONT KNOW WHAT IM DOING WRONG